OpenStudy (anonymous):

The sum of two numbers is37 . The smaller number is 7 less than the larger number. What are the numbers?

OpenStudy (anonymous):

first, define your variables: let your smaller # = n - 7 let your larger # = n

OpenStudy (anonymous):

then add them together since the sum of the two numbers is 37: n - 7 + n = 37

OpenStudy (anonymous):

2n - 7 = 37

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

you're welcome!